# Runge-Kutta Approximations for Direct Coning Compensation Applying Lie Theory

### v1

@@
Abstract
The integration of gyroscope measurements is an essential task for most navigation systems. Modern vehicles typically use strapdown systems, such that gyro integration requires coning compensation to account for the sensor’s rotation during the integration. Many coning compensation algorithms have been developed and a few are reviewed. This work introduces a new class of coning correction algorithm built directly from the classical Runge-Kutta integration routines. A simple case is shown to collapse to one of the most popular coning algorithms and a clear procedure for generating higher-order algorithms is presented.
@@ 1 Introduction
Strapdown inertial measurement units (IMUs) are ubiquitous sensors in modern aerospace systems, automobiles, ground robotics, consumer electronics, and other platforms. Typical IMUs are composite sensors consisting of both gyroscopes (gyros) and accelerometers. In their simplest configuration, a three-axis IMU consists of three orthogonally arranged gyros and three orthogonally arranged accelerometers. Each gyro within the IMU measures the instrument’s inertial angular velocity in a particular instrument-fixed direction. Likewise, each accelerometer measures the instrument’s sensed acceleration (specific force). This work focuses primarily on the integration of the gyro measurements to obtain an estimate of the inertial change in IMU orientation.
The conventional approach for integrating gyro measurements follows a multi-frequency scheme. This scheme, which has been written about extensively elsewhere [1, 2], consists of three important time intervals (Fig. 1). The shortest interval is the sensor interval (sometimes called the “subminor interval”), corresponding to the time over which the instrument integrates the raw gyro signal to produce a discrete measurement for processing. These discrete measurements are produced at a high rate, generally on the order of several kHz. The middle interval is the minor interval in which the kHz measurements are integrated to produce a rotation measurement suitable for consumption by a navigation filter. This is usually done within the IMU and the minor interval corresponds to the time between measurements of the IMU data stream seen by a user. Finally, the navigation interval (sometimes called the “major interval”) is the time between state estimates within the navigation filter.
@@
The raw signal from each gyro is typically integrated in hardware on a per-channel basis. The independent integration of gyro channels introduces error in a rotating strapdown IMU since the gyros change orientation over the course of integration and rotations are noncommutative. We refer to the error introduced by this process as coning error and attempt to correct for this through coning compensation algorithms. Many older IMUs also integrate gyro measurements on a per-channel basis over the minor interval due to computational constraints in keeping up with a kHz data stream and then perform higher-order coning corrections using a multitude of sensor-interval integrations [2, 3]. Although many IMUs continue to separate the commutative and noncommutative integrations at the minor interval, this no longer necessary with modern computational capabilities. Nevertheless, coning correction remains necessary over the sensor interval whenever raw gyro signals are integrated (or averaged) on a per-channel basis.
The main contribution of this work is generalizing coning compensation algorithms to higher-order integrators. Attitude propagation is traditionally stated as an initial value problem where the ordinary differential equation (ODE) is referred to as the Bortz equation. To our knowledge, we are the first to equate the Bortz equation to the inverse-Jacobian of the Lie algebra. Simplifying the ODE, motivated by Lie theory, enables both usage of traditional Runge-Kutta integrators and approximate solutions discarding higher-order terms. The Lie-theoretic perspective highlights a need to model angular rates from integrated measurements. Our second contribution is fitting polynomial models for instantaneous rates from integrated measurements. Aggregating multiple measurements enables higher-order polynomial models which in-turn enables higher-order integrator methods. Our final contribution is quantifying variations in integration error based on solver order, polynomial order of rate models, and approximate solutions to the ODE. We provide new methods for propagating attitude for inertial navigation leveraging consecutive integrated rate measurements. Our results can be used to reduce integration error or as method to increase step size (potentially reducing computations) while preserving fidelity.
@@ 2 Attitude Representations and Kinematics
Consider an IMU with body frame $B$ that is rotating with respect to an inertial frame $I$ at an angular velocity of $\boldsymbol{\omega}$. Our notation assumes that $\boldsymbol{\omega}$ is expressed in the rotating frame $B$. We may write the rotation between frames $I$ and $B$ using the direction cosine matrix (DCM) T, such that a vector x is transformed according to
@@
where $\emph{{x}}_{I}$ is the vector expressed in frame $I$ and $\emph{{x}}_{B}$ is the vector expressed in frame $B$. Here, we assume the passive interpretation of attitude such that T describes a change of basis [4, 5] transforming a vector expressed in frame $I$ to that same vector expressed in frame $B$. The DCM $\emph{{S}}=\emph{{T}}^{-1}$ describes the reverse transfomration.
Every DCM T (or S) belongs to the group $SO(3)$ consisting of $3\times 3$ proper orthogonal matrices. As a consequence, every DCM $\emph{{T}}\in SO(3)$ has three free parameters and six constraints. Although the DCM is convenient for performing frame transformations of vectors, the constraints make perfectly on-manifold integration of the DCM more difficult. It is preferable to perform the integration on a minimal attitude representation (i.e., a three-parameter representation). Many such representations exist [4], but we will only consider the rotation vector here.
We may interpret the problem of attitude representation, kinematics, and integration through either a conventional aerospace framework or from a Lie-theoretic framework. These interpretations are equivalent and lead to exactly the same equations. However, as we will see, each also comes with its own history, vocabulary, and observations.
@@ 2.1 Conventional Aerospace Framework
Euler’s theorem states that any arbitrary 3-axis rotation may be described by the rotation about a single axis e and a single angle $\varphi$. Taking the product of these forms the rotation vector $\boldsymbol{\phi}$, which is a $3\times 1$ vector pointing in the direction of e with a magnitude of $\varphi$. Thus, we write the rotation vector as
@@
where $\boldsymbol{\omega}$ is the angular velocity of frame $B$ with respect to frame $I$ as expressed in frame $B$. We sometimes refer to eq. (3) as the “Bortz equation” (especially in the IMU literature) after an influential derivation presented by the American engineer John Bortz in 1971 [6]. A somewhat simpler way of arriving at the same result using quaternions may be found in Ref. [4].
The elements e and $\varphi$ of the rotation vector are related to the DCM through the finite rotation formula [4]
@@
We may also write the DCM as the matrix exponential
@@
though most navigators compute the DCM in practice using eq. (4) for computational reasons. We also recognize the DCM kinematics to be
@@
such that $\dot{\emph{{T}}}^{T}=\dot{\emph{{S}}}$. All of these notions are routinely encountered in aerospace textbooks [4, 7]. The reader is reminded that $\boldsymbol{\omega}\neq\dot{\boldsymbol{\phi}}$, in general, with an important exception of $\boldsymbol{\phi}=\textbf{0}_{3\times 1}$. These things will be important later.
@@ 2.2 Lie-Theoretic Framework
One of the primary advantages of the Lie-theoretic approach is that an exponential map can be used to relate a Lie algebra with a corresponding Lie group. In the present case were are interested in the Lie group of $SO(3)$. The corresponding Lie algebra $so(3)$ may be constructed with the wedge operator $(\,\cdot\,)^{\wedge}:\mathbb{R}^{3}\rightarrow so(3)$. The exponential map $\exp:so(3)\rightarrow SO(3)$ relating $\emph{{T}}\in SO(3)$ and $\emph{{s}}^{\wedge}\in so(3)$ is simply the matrix exponential
@@
when speaking of $so(3)$. This is well-known [8, 9]. We emphasize, however, that group transformations extend elegantly to higher-dimensional spaces and other groups with relevance to navigation and robotics. For example, the groups $SO(4)$ and $SE(3)$ are both inherently 6-dimensional with corresponding Lie Algebras $so(4)$ and $se(3)$, respectively. See Ref. [10] for a concise generalized treatment.
The kinematics from eq. (7a) can be restated for S using Lie-theoretic notation [11, 10]
@@
Since $\emph{{T}}^{-1}=\emph{{T}}^{\top}\in SO(3)$, one implication of eq. (10) is that $\emph{{T}}\dot{\emph{{T}}}^{\top}=\boldsymbol{\omega}^{\wedge}\in so(3)$. As a reciprocal operation to $(\,\cdot\,)^{\wedge}$, the “vee operator” converts skew-symmetric matrices to their vector representations $(\,\cdot\,)^{\vee}:so(3)\rightarrow\mathbb{R}^{3}$}. We now can express $\emph{{T}}\dot{\emph{{T}}}^{\top}$ as a vector decomposed as a matrix-vector product
@@
Since the Jacobian is invertable for all s [10], the differential equation in the Lie algebra is simply
@@
For $SO(3)$, the inverse right-Jacobian $\emph{{J}}_{\boldsymbol{\phi}}^{-1}$ has the closed form (c.f. equation (10.86) [10], also [12, 13, 14, 15]) [^6] Note that the sign of the middle term $(1/2)\boldsymbol{\phi}^{\wedge}$ may vary from one reference to another depending on the choice of an active or passive representation of attitude.
$\emph{{J}}_{\emph{{s}}}^{-1}=\emph{{I}}+\frac{1}{2}\boldsymbol{\phi}^{\wedge}+\left(\frac{1}{\varphi^{2}}-\frac{1+\cos\varphi}{2\varphi\sin\varphi}\right)\left(\boldsymbol{\phi}^{\wedge}\right)^{2}$ (14)
It is helpful to observe
@@
such that to third order in $\varphi=\|\boldsymbol{\phi}\|$ the right-Jacobian is sometimes approximated as
$\emph{{J}}_{\emph{{s}}}^{-1}\approx\emph{{I}}+\frac{1}{2}\boldsymbol{\phi}^{\wedge}+\frac{1}{12}\left(\boldsymbol{\phi}^{\wedge}\right)^{2}$ (16)
@@
Recalling that $\boldsymbol{\phi}^{\wedge}=[\boldsymbol{\phi}\times]$ produces the desired equivalence.
@@ 3 Conventional Coning Correction
We may compute $\emph{{T}}(\Delta\boldsymbol{\phi}_{k})$ using the either the finite rotation formula from eq. (4) or the exponential map from Eqs. (6a) or (8). In practice we typically use the former, though it does not matter much one way or the other.
@@
What makes this challenging in practice is that the gyros within an IMU do not directly measure either $\Delta\boldsymbol{\phi}_{k}$ or $\dot{\boldsymbol{\phi}}$. Instead, in an ideal IMU with perfectly orthogonal gyros, each gyro channel measures $\omega_{1}$, $\omega_{2}$, or $\omega_{3}$ in the rotating body frame. We assemble these into the angular velocity vector $\boldsymbol{\omega}=[\omega_{1};\omega_{2};\omega_{3}]$. Thus, per-channel integration of the angular velocity produced by a typical rate-integrating gyro (e.g., a ring laser gyro) actually produces a measurement of
@@
The objective is to write what we want ($\Delta\boldsymbol{\phi}_{k}$) in terms of what we can actually measure ($\Delta\boldsymbol{\theta}_{k}$). The difference in the two is principally due to the noncommutitavity terms in the Bortz equation [the right-most term in eq. (3)]. Thus, integrating the Bortz equation to obtain $\boldsymbol{\phi}$ produces
@@
We refer to $\boldsymbol{\beta}$ as the coning correction term defined as
@@
Thus, as is clear from eq. (23), the coning term $\boldsymbol{\beta}$ is a compensation to the measured $\boldsymbol{\theta}$ that attempts to account for the changing orientation of frame $B$ over the integration. Clearly, for a one dimensional problem where $\boldsymbol{\omega}$ remains fixed in the body frame, we find that $\boldsymbol{\beta}_{k}=\textbf{0}$ since $\boldsymbol{\phi}\times\boldsymbol{\omega}=\textbf{0}$.

### v2

@@
Runge-Kutta Approximations for Direct Coning Compensation Applying Lie Theory
John A. Christian  [^1] Associate Professor, Guggenheim School of Aerospace Engineering, Associate Fellow AIAA. Michael R. Walker II [^2] Principal Member Technical Staff [^3] Corresponding author, mrw2@ieee.org Wyatt Bridgman [^4] Now, Senior Data Scientist with C3 AI, Falls Church, VA Michael J. Sparapany [^5] Principal Member Technical Staff
@@ 1 Introduction
Strapdown inertial measurement units (IMUs) are ubiquitous sensors in modern aerospace systems, automobiles, ground robotics, consumer electronics, and other platforms. Typical IMUs are composite sensors consisting of both gyroscopes (gyros) and accelerometers. In their simplest configuration, a three-axis IMU consists of three orthogonally arranged gyros and three orthogonally arranged accelerometers. Each gyro within the IMU measures the instrument’s inertial angular velocity in a particular instrument-fixed direction. Likewise, each accelerometer measures the instrument’s sensed acceleration (specific force). This work focuses primarily on the integration of the gyro measurements to obtain an estimate of the inertial change in IMU orientation. Such IMU integration is a central component to the propagate function within most aerospace navigation filters [Markley:2014, crassidisThreeAxisAttitudeEstimation2016a, carpenterNavigationFilterBest2025, christianFundamentalsSpacecraftOptical2026].
The conventional approach for integrating gyro measurements follows a multi-frequency scheme. This scheme, which has been written about extensively elsewhere [Ignagni:1990, Savage:1998], consists of three important time intervals (Fig. 1). The shortest interval is the sensor interval (sometimes called the “subminor interval”), corresponding to the time over which the instrument integrates the raw gyro signal to produce a discrete measurement for processing. These discrete measurements are produced at a high rate, generally on the order of several kHz. The middle interval is the minor interval in which the kHz measurements are integrated to produce a rotation measurement suitable for consumption by a navigation filter. This is usually done within the IMU and the minor interval corresponds to the time between measurements of the IMU data stream seen by a user. Finally, the navigation interval (sometimes called the “major interval”) is the time between state estimates within the navigation filter.
@@
The raw signal from each gyro is typically integrated in hardware on a per-channel basis. The independent integration of gyro channels introduces error in a rotating strapdown IMU since the gyros change orientation over the course of integration and rotations are noncommutative. Error introduced by this process is referred to as coning error and corrections are attempted through coning compensation algorithms[Ignagni:1996, Savage:1998]. Many older IMUs also integrate gyro measurements on a per-channel basis over the minor interval due to computational constraints in keeping up with a kHz data stream and then perform higher-order coning corrections using a multitude of sensor-interval integrations [Savage:1998, Ignagni:1996]. Although many IMUs continue to separate the commutative and noncommutative integrations at the minor interval, this no longer necessary with modern computational capabilities. Nevertheless, coning correction remains necessary over the sensor interval whenever raw gyro signals are integrated (or averaged) on a per-channel basis.
The main contribution of this work is generalizing coning compensation algorithms to higher-order integrators. Gyros do not provide absolute measures of vehicle attitude. Instead, vehicle attitude states must be propagated - or integrated - using inertial measurements. Attitude propagation fundamentally represents an initial value problem where the ordinary differential equation (ODE) depends on the coordinate choice for state variables. Quaternions are a common choice for attitude representation within the navigation community, and high-order integration methods have been demonstrated for instantaneous rate measurements [andrleGeometricIntegrationQuaternions2013a]. Low-order methods have been demonstrated for rate-integrating gyro [crassidisThreeAxisAttitudeEstimation2016a]. For attitude represented in exponential coordinates (also known as the rotation vector), the ODE is referred to as the Bortz equation in the navigation community. This work makes explicit the equivalence between the Bortz equation and the inverse-Jacobian of the Lie algebra. Recognizing the Bortz equation as an invertible matrix simplifies the state-propagation ODE. The simplified ODE conveniently fits into traditional Runge-Kutta integrators, and approximate solutions are available by discarding higher-order terms. The Lie-theoretic perspective highlights a need to model angular rates from integrated measurements. The second contribution is fitting polynomial models for instantaneous rates from integrated measurements. Aggregating multiple measurements enables higher-order polynomial models which in-turn enables higher-order integrator methods. The third contribution is quantifying variations in integration error based on solver order, polynomial order of rate models, and approximate solutions to the ODE. New methods for propagating attitude are provided for inertial navigation leveraging consecutive integrated rate measurements. The combination of curve-fitting and ODE solver selection decouples solver order from the number of integrated measurements. Results can be used to reduce integration error or as a method to increase step size (potentially reducing computations) while preserving fidelity.
@@ 2 Attitude Representations and Kinematics
Consider an IMU with body frame $B$ that is rotating with respect to an inertial frame $I$ at an angular velocity of $\boldsymbol{\omega}$. This notation assumes that $\boldsymbol{\omega}$ is expressed in the rotating frame $B$. The rotation between frames $I$ and $B$ may be written using the direction cosine matrix (DCM) T, such that a vector x is transformed according to
@@
where $\emph{{x}}_{I}$ is the vector expressed in frame $I$ and $\emph{{x}}_{B}$ is the vector expressed in frame $B$. Here, the passive interpretation of attitude is assumed such that T describes a change of basis [Markley:2014, Zanetti:2019] transforming a vector expressed in frame $I$ to that same vector expressed in frame $B$. The DCM $\emph{{S}}=\emph{{T}}^{-1}$ describes the reverse transformation.
Every DCM T (or S) belongs to the group $SO(3)$ consisting of $3\times 3$ proper orthogonal matrices. As a consequence, every DCM $\emph{{T}}\in SO(3)$ has three free parameters and six constraints. Although the DCM is convenient for performing frame transformations of vectors, the constraints make perfectly on-manifold integration of the DCM more difficult. It is preferable to perform the integration on a minimal attitude representation (i.e., a three-parameter representation). Many such representations exist [Markley:2014], but only the rotation vector is considered here for brevity.
The problem of attitude representation, kinematics, and integration may be interpreted through either a conventional aerospace framework or from a Lie-theoretic framework. These interpretations are equivalent and lead to exactly the same equations. However, as will be shown, each also comes with its own history, vocabulary, and observations.
@@ 2.1 Conventional Aerospace Framework
Euler’s rotation theorem states that any arbitrary 3-axis rotation may be described by the rotation about a single axis e and a single angle $\varphi$ [eulerFormulaeGeneralesPro1776, Markley:2014, christianFundamentalsSpacecraftOptical2026]. Taking the product of these forms the rotation vector $\boldsymbol{\phi}$, which is a $3\times 1$ vector pointing in the direction of e with a magnitude of $\varphi$. Thus, the rotation vector is written as
@@
where $\boldsymbol{\omega}$ is the angular velocity of frame $B$ with respect to frame $I$ as expressed in frame $B$. Since $\boldsymbol{\phi}$ is a minimal parameterization, singularities exist. However, for the exponential coordinates singularities are not encountered for $\varphi<2\pi$} [stuelpnagelParametrizationThreeDimensionalRotation1964].
Equation (3) is sometimes referred to as the “Bortz equation” (especially in the IMU literature) after an influential derivation presented by the American engineer John Bortz in 1971 [Bortz:1971]. A somewhat simpler way of arriving at the same result using quaternions may be found in Ref. [Markley:2014].
The elements e and $\varphi$ of the rotation vector are related to the DCM through the finite rotation formula [Markley:2014]
@@
The DCM may also be written as the matrix exponential
@@
though most navigators compute the DCM in practice using eq. (4) for computational reasons. The DCM kinematics are also recognized to be
@@
such that $\dot{\emph{{T}}}^{\top}=\dot{\emph{{S}}}$. All of these notions are routinely encountered in aerospace textbooks [Markley:2014, Crassidis:2012]. The reader is reminded that $\boldsymbol{\omega}\neq\dot{\boldsymbol{\phi}}$, in general, with an important exception of $\boldsymbol{\phi}=\textbf{0}_{3\times 1}$. These things will be important later.
@@ 2.2 Lie-Theoretic Framework
One of the primary advantages of the Lie-theoretic approach is that an exponential map can be used to relate a Lie algebra with a corresponding Lie group. Although extensions to other Lie groups are possible, this work focuses strictly on $SO(3)$. The corresponding Lie algebra $so(3)$ may be constructed with the wedge operator $(\,\cdot\,)^{\wedge}:\mathbb{R}^{3}\rightarrow so(3)$. The exponential map $\exp:so(3)\rightarrow SO(3)$ relating $\emph{{T}}\in SO(3)$ and $\emph{{s}}^{\wedge}\in so(3)$ is simply the matrix exponential
@@
when speaking of $so(3)$. This is well-known [Wang:2008, Forster:2017]. It is worth emphasizing, however, that group transformations extend elegantly to higher-dimensional spaces and other groups with relevance to navigation and robotics. For example, unit-quaternions compose a Lie Group [andrleGeometricIntegrationQuaternions2013a]. The groups $SO(4)$ and $SE(3)$ are both inherently 6-dimensional with corresponding Lie Algebras $so(4)$ and $se(3)$, respectively. See Ref. [chirikjian_stochastic_2012] for a concise generalized treatment.
The kinematics from eq. (7a) can be restated for S using Lie-theoretic notation [iserles2000lie, chirikjian_stochastic_2012]
@@
Since $\emph{{T}}^{-1}=\emph{{T}}^{\top}\in SO(3)$, one implication of eq. (10) is that $\emph{{T}}\dot{\emph{{T}}}^{\top}=\boldsymbol{\omega}^{\wedge}\in so(3)$. As a reciprocal operation to $(\,\cdot\,)^{\wedge}$, the “vee operator” converts skew-symmetric matrices to their vector representations $(\,\cdot\,)^{\vee}:so(3)\rightarrow\mathbb{R}^{3}$}. Now $\emph{{T}}\dot{\emph{{T}}}^{\top}$ can be expressed as a vector decomposed as a matrix-vector product
@@
Since the Jacobian is invertible for all s [chirikjian_stochastic_2012], the differential equation in the Lie algebra is simply
@@
For $SO(3)$, the inverse right-Jacobian $\emph{{J}}_{\boldsymbol{\phi}}^{-1}$ has the closed form (c.f. equation (10.86) [chirikjian_stochastic_2012], also [hausdorff1906symbolische, iserles1984solving, Muller:2023, sola2018micro]) [^6] Note that the sign of the middle term $(1/2)\boldsymbol{\phi}^{\wedge}$ may vary from one reference to another depending on the choice of an active or passive representation of attitude.
$\emph{{J}}_{\boldsymbol{\phi}}^{-1}=\emph{{I}}+\frac{1}{2}\boldsymbol{\phi}^{\wedge}+\left(\frac{1}{\varphi^{2}}-\frac{1+\cos\varphi}{2\varphi\sin\varphi}\right)\left(\boldsymbol{\phi}^{\wedge}\right)^{2}$ (14)
It is helpful to observe [Savage:1998]
@@
such that to third order in $\varphi=\|\boldsymbol{\phi}\|$ the right-Jacobian may be approximated as
$\emph{{J}}_{\boldsymbol{\phi}}^{-1}\approx\emph{{I}}+\frac{1}{2}\boldsymbol{\phi}^{\wedge}+\frac{1}{12}\left(\boldsymbol{\phi}^{\wedge}\right)^{2}$ (16)
@@
which is equivalent to (3). Recalling that $\boldsymbol{\phi}^{\wedge}=[\boldsymbol{\phi}\times]$, and using $\boldsymbol{\phi}\times\left[\boldsymbol{\phi}\times\boldsymbol{\omega}\right]=\left(\boldsymbol{\phi}^{\wedge}\right)^{2}\boldsymbol{\omega}$}, produces the desired equivalence.
For completeness,
$J_{\boldsymbol{\phi}}=\emph{{I}}-\frac{1-\cos\varphi}{\varphi^{2}}\boldsymbol{\phi}^{\wedge}+\frac{\varphi-\sin\varphi}{\varphi^{3}}\left(\boldsymbol{\phi}^{\wedge}\right)^{2}$ (18)
using the right-Jacobian convention [chirikjian_stochastic_2012].
The limit of (15) as $\varphi\rightarrow 0$} is simply $1/12$. Consequently, the Jacobian is well behaved over the domain $\boldsymbol{\phi}<2\pi$} when $\boldsymbol{\phi}$ is expressed in exponential coordinates. For other three-dimensional parameterizations of rotations (e.g., Euler angles), singularity are encountered in similar domains [stuelpnagelParametrizationThreeDimensionalRotation1964].
@@ 3 Conventional Coning Correction
Computation of $\emph{{T}}(\Delta\boldsymbol{\phi}_{k})$ may use either the finite rotation formula from eq. (4) or the exponential map from eqs. (6a) or (8). In practice the former is arguably more prevalent, though it does not matter much one way or the other.
@@
What makes this challenging in practice is that the gyros within an IMU do not directly measure either $\Delta\boldsymbol{\phi}_{k}$ or $\dot{\boldsymbol{\phi}}$. Instead, in an ideal IMU with perfectly orthogonal gyros, each gyro channel measures $\omega_{1}$, $\omega_{2}$, or $\omega_{3}$ in the rotating body frame. The measurements compose the angular velocity vector $\boldsymbol{\omega}=[\omega_{1};\omega_{2};\omega_{3}]$. Thus, per-channel integration of the angular velocity produced by a typical rate-integrating gyro (e.g., a ring laser gyro) actually produces a measurement of
@@
The objective is to write the desired quantity ($\Delta\boldsymbol{\phi}_{k}$) in terms of available measurements ($\Delta\boldsymbol{\theta}_{k}$). The difference in the two is principally due to the noncommutitavity terms in the Bortz equation [the right-most term in eq. (3)]. Thus, integrating the Bortz equation to obtain $\boldsymbol{\phi}$ produces
@@
The coning correction term $\boldsymbol{\beta}$ is defined as
@@
Thus, as is clear from eq. (24), the coning term $\boldsymbol{\beta}$ is a compensation to the measured $\boldsymbol{\theta}$ that attempts to account for the changing orientation of frame $B$ over the integration. Clearly, for a one dimensional problem where $\boldsymbol{\omega}$ remains fixed in the body frame, it is found that $\boldsymbol{\beta}_{k}=\textbf{0}$ since $\boldsymbol{\phi}\times\boldsymbol{\omega}=\textbf{0}$.